You are here

Beam delivery update

header

10 December 2014

Things are running relatively well, with the cyclotron delivering 9.5 µA to BL2A, ~95 µA to BL2C4 for isotope production, and beam as needed to BL1B for PIF. BL2A is running at a slightly reduced current to give a little more headroom before hitting the 10 µA licence limit for UCx targets because of changes in the cyclotron tune with BL1B coming on an off. BL2A may go off for a few hours if RIB is not needed overnight, but otherwise we expect to continue like this for the next few days.

At ISAC, stable-beam tuning for FrPNC experiment S1218 has been completed and 211Fr is available; delivery is on hold while the experimenters continue to set things up. That experiment is scheduled to run through the weekend. Stable-beam delivery from OLIS has finished for the year – today’s detector tests at DRAGON wrapped up late this afternoon – and OLIS will be used for operator training before it too is shut down.

-- ACM